14. 4:6::2:X

complete the ratio. X =______

15. A patient who weighs 168 lb is prescribed 15 mg of theophylline per kilogram of body weight intravenously as a loading dose. The drug is available at 250 mg/10 mL. How many milliliters should you prepare for this patient? Round to the nearest 10th of a mL. ___________mL

19. What technique should you use when measuring a liquid drug in a medicine cup to ensure an accurate dose?

a. Use a syringe to draw up the drug and then place it in the medicine cup.

b. Hold the cup above eye level while pouring the drug.

c. Hold the cup below eye level while pouring the drug.

d. Place the cup on a table and bend to look straight at the mark
Answer #1

14. Answer.

4:6::2:X

It means,

4/6 = 2/X

Now cross multiply,

4X = 12

X= 12/4

X= 3

The value of X is 3.

15. Answer.

Order

Theophylline.

Weight of the patient is = 168 lb = 76.2035 kgs.

~ 76.2 kgs.

Needed dose 15 mg/kg of body weight.

= 15 * 76.2 kg ( body weight)

= 1143 mg total drug is needed.

Desired dose= 1143 mg

Dose in hand = 250 mg.

quantity of dilution = 10 ml.

the volume of drug needed =

Desired dose / Dose in hand * quantity.

= 1143 mg / 250 mg * 10 ml.

= 45.72 ml.  

~46 ml.

​​​​​​19. Answer.

A medicine cup is used to measure liquids or fluids, also to measure drugs in liquid form, this cup has a measuring scale made on it which makes the measurement of liquid medication easy.

Option (B) is the right answer.   

Hold the cup above eye level while pouring the drug.

By doing this the exact quantity of medication in the cup can be easily seen also it facilitates easier withdrawal of medication.

option A) is the wrong answer because there is no need to measure the drug with the syringe as the cup has already measuring scale on it.

Option( C) and option( D ) are wrong answers because while pouring the medication by keeping the cup below the eye level may not facilitate correct measurement and also it becomes hard to pour the exact amount of medicine into the cup.

Putting the glass on the table while pouring medicines cause refraction of light, the light wave bends when it passes through the cup and makes it harder to know the exact quantity.

​​.